OpenStudy (anonymous):

Answer Please ? 2x 4 - 20x 2 - 78

OpenStudy (anonymous):

PEMDAS RULE

OpenStudy (anonymous):

8-40-78

OpenStudy (mathstudent55):

Are you trying to factor?

OpenStudy (anonymous):

i think it's an expression

OpenStudy (mathstudent55):

@Jaydee Is this a factoring problem: Factor the trinomial \(2x^4 - 20x^2 - 78\) Or is this an expression that needs to be evaluated: Evaluate: \(2 \times 4 - 20 \times 2 - 78 \) \( \LARGE ? \) Or perhaps something else?

OpenStudy (anonymous):

he's not even trying -_-

OpenStudy (anonymous):

8-40-78 -32-78 =-110